OpenStudy (anonymous):

The quadratic x^2 + 5x – 14 has what type of roots?

OpenStudy (campbell_st):

use the descriminant\[b^2 - 4ac\] =0 equal roots >0 real, <0 complex in your question 5^2 - 4 x 1 x (-14) = 81 so real roots that are rational since 81 is a square number

OpenStudy (anonymous):

so one rational double or two rational?

OpenStudy (campbell_st):

all quadratics have a max of 2 roots.... if \[b^2 -4ac >0\] the roots are unequal real, (x +7)(x-2) means x = -7, 2 these are the roots